Fwd: EPI2 status update: dome/ weekend shutdown

maurice stevens maurice at stanford.edu
Wed Nov 10 15:48:02 PST 2010












-------- Original Message --------
Subject: 	EPI2 status
Date: 	Tue, 09 Nov 2010 16:52:11 -0800
From: 	maurice <maurice at stanford.edu>
To: 	epi2 at snf.stanford.edu



After may etches, the lower temp is reading 883 c but the upper is only 
reading 865 (both should be reading 885c).  So it looks like the lower 
dome cleared but we can see the upper dome is still coated.
The system is purging with N2 overnight and we'll pull the upper dome 
tomorrow.
After the dome is replaced it usually takes a day or so before we can 
get haze free deposition.

Also, facilities is planning to shut off the process cooling water in 
CISX.  The process cooling water for EPI2 is supplied from CISX...so 
EPI2 will be down this weekend.


-m












It looks  like the lower dome got coated.

I ran the 20KW temp test and the lower zone was reading 870c (should be 
885)  The upper zone was reading 849 (should be 885).

After running some long constant power/ low slitH2 etches the lower is 
879 and the upper 856.

I'll continue etching and running temp tests tomorrow.

-m
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/epi2/attachments/20101110/5ade59b8/attachment.html>


More information about the epi2 mailing list